# ReZero: Boosting MCTS-based Algorithms by Backward-view and Entire-buffer Reanalyze

### v4

@@ Abstract
Monte Carlo Tree Search (MCTS)-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains. These algorithms employ the reanalyze process to enhance sample efficiency from stale data, albeit at the expense of significant wall-clock time consumption. To address this issue, we propose a general approach named ReZero to boost tree search operations for MCTS-based algorithms. Specifically, drawing inspiration from the one-armed bandit model, we reanalyze training samples through a backward-view reuse technique which obtains the value estimation of a certain child node in advance. To further adapt to this design, we periodically reanalyze the entire buffer instead of frequently reanalyzing the mini-batch. The synergy of these two designs can significantly reduce the search cost and meanwhile guarantee or even improve performance, simplifying both data collecting and reanalyzing. Experiments conducted on Atari environments and board games demonstrate that ReZero substantially improves training speed while maintaining high sample efficiency. The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
@@ 1 Introduction
As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) [1] has acquired achievements and applications across diverse fields, including interactive gaming [2], autonomous vehicles [3], and natural language processing [4]. Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency. These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real-world scenarios. In response to this problem, numerous model-based RL [5, 6] methods have emerged. These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
@@
Within this field, Monte Carlo Tree Search [7] (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm [8] into the tree search process and has achieved promising results in in a wide range of scenarios. Specifically, AlphaZero [9] plays a big role in combining deep reinforcement learning with MCTS, achieving notable accomplishments that can beat top-level human players. While it can only be applied to environments with perfect simulators, MuZero [10] extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces [11], environments with stochastic dynamics [12], offline RL training scenarios [13], and etc. All these MCTS-based algorithms [9, 10, 14] made valuable contributions to the universal applicability of the MCTS+RL paradigm.
@@
However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state. Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments [15], further amplifying the speed disadvantage. On the other hand, during the reanalyze [13] process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on the training mini-batch. The time required to process samples thus increases as a trade-off for high sample efficiency. The excessive time cost has become a bottleneck hindering the further promotion of these algorithms.
@@
Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms. On the one hand, SpeedyZero [16] greatly diminishes algorithms’ time overheads by deploying a parallel execution training pipeline; however, it demands additional computational resources. On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra demands. PTSAZero [17], for instance, compresses the search space via state abstraction, decreasing the time cost per search. Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times (the number of simulations) to assure their performance. From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible.
@@
In this paper, we introduce ReZero, a new approach designed to boost the MCTS-based algorithms. Firstly, inspired by the one-armed bandit model [18], we propose a backward-view reanalyze technique that proceeds in the reverse direction of the trajectories, utilizing previously searched root values to bypass the exploration of specific child nodes, thereby saving time. We present several propositions to articulate our insights and analyzed the convergence of the algorithm based on the bandit model. Secondly, to better adapt to our proposed backward-view reanalyze technique, we have devised a novel pipeline that concentrates MCTS calls within the reanalyze process and periodically reanalyzes the entire buffer after a fixed number of training iterations. This entire-buffer reanalyze not only reduces the number of MCTS calls but also better leverages the speed advantages of parallelization. Our approach can be easily applied to the MCTS-based algorithm family and synergistically implemented alongside existing acceleration methods. In addition, it will not bring about any overhead in computation resources. Empirical experiments (Section 5) show that our approach yields good results in both single-agent environment (Atari [19]) and two-player board games [14], greatly improving the training speed while maintaining or even improving sample efficiency. Ablation experiments explore the impact of different reanalyze frequencies and the acceleration effect of backward-view reuse on a single search. The main contributions of this paper can be summarized as follows:
@@
  • We design a method to speed up a single search by the backward-view reanalyze technique. Theoretical support for the effectiveness of our proposed method is also provided.
@@
  • We propose a simplified framework with the entire-buffer reanalyze mechanims that further reduces the number of MCTS calls and enhance its parallelization, thus boosting MCTS-based algorithms.
@@
  • We conduct several experiments on diverse environments and investigate ReZero through ablations.
@@ 2.1 MCTS-based Algorithms
AlphaGo [14] and AlphaZero [9] combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero [10] combines tree search with a learned value equivalent model [20], successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. EfficientZero [21] further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Another works [11, 12] extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. MuZero Unplugged [13] applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Gumbel MuZero [22] changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@ 2.2 MCTS Acceleration
Recent research has focused on accelerating MCTS-based algorithms. The work [16] reduces the algorithm’s time overhead by designing a parallel system; however, this method requires more computational resources and involves some adjustments for large batch training. PTSAZero [17] narrows the search space through state abstraction, which amalgamates redundant information, thereby reducing the time cost per search. KataGo [23] use a naive information reuse trick. They save sub-trees of the search tree and serve as initialization for the next search. However, our proposed backward-view reanalyze technique is fundamentally different from this naive forward-view reuse and can enhance search results while saving time. To our knowledge, we are the first to enhance MCTS by reusing information in a backward-view. Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which [16, 17, 22, 24] are orthogonal to our contributions.
@@ 3.1 MuZero
MuZero [10] is a fundamental model-based RL algorithm that incorporates a value-equivalent model [25] and leverages MCTS for planning within a learned latent space. The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
@@
The representation model transforms a last observation sequences $o_{t-l:t}$ into a corresponding latent state $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value $v_{t}$. These outputs are instrumental in guiding the agent’s action selection throughout its MCTS. Lastly the agent samples the best action $a_{t}$ following the searched visit count distribution. MuZero Reanalyze is an advanced version of the original MuZero algorithm. This variant enhances the model’s accuracy by conducting a fresh MCTS on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection. Traditional uses of the algorithm have intertwined reanalysis with training, while we suggests a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology.
@@ 3.2 Bandit-view Tree Search
A stochastic bandit has $K$ arms, and playing each arm means sampling a reward from the corresponding distribution. For a search tree in MCTS, the root node can be seen as a bandit, with each child node as an arm. The left side of Figure 1 illustrates this idea. However, as the policy is continuously improved during the search process, the reward distribution for the arms should change over time. Therefore, UCT [27] modeled the root node as a non-stationary stochastic bandit with a drift condition:
$\mathbb{P}(\overline{X}_{is}-\mu_{i}\geq\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})\>\text{and}\>\mathbb{P}(\overline{X}_{is}-\mu_{i}\leq-\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})$ (2)
Where $\overline{X}_{is}$ is the average reward of the first $s$ samples of arm $i$. $\mu_{i}$ is the limit of $\mathbb{E}[\overline{X}_{is}]$ as $s$ approaches infinity, which indicates that the expectation of the node value converges. $C$ is an appropriate constant characterizes the rate of concentration.
Based on this modeling, UCT uses the bandit algorithm UCB1 [8] to select child nodes. AlphaZero [9] inherits this concept and employs a variant formula:
@@
where $s$ is the state corresponding to the current node, $a$ is the action corresponding to a child node. $Q(s,a)$ is the mean return of choosing action $a$, $P(s,a)$ is the prior score of action $a$, $N(s,a)$ is the total time that $a$ has been chosen, ${\sum}_{b}N(s,b)$ is the total time that $s$ has been visited.
@@ 4.1 Backward-view Reanalyze
Our algorithm design stems from a simple inspiration: if we could know the true state-value of a child node in advance, we could save the search for it, thus conserving search time. As shown on the right side of Figure 1, we directly use the true expected value $\mu_{A}$ to evaluate the quality of Arm A, thereby eliminating the need for the back-propagated value to calculate the score in 3. This results in the process occurring within the gray box being omitted. Indeed, this situation can be well modeled as an one-armed bandit[18], which also facilitates our subsequent theoretical analysis.
@@
Figure 1: The connection between MCTS and bandits. Left shows the tree search in a bandit-view. When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$. For the root node, the traversal, evaluation and back-propagation occurring in the sub-tree can be approximated as sampling from an non-stationary distribution. Thus the root node can be seen as a non-stationary bandit. Right shows the one-armed bandit case. Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
@@
Driven by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance. However, the true expectation is always unknown. Consequently, we resort to using the root value obtained from MCTS as an approximate substitute. Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$  [^1] The subscript denotes the trajectory, the superscript denotes the time step in the trajectory. in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$. Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$. However, there is a temporal contradiction. $S_{0}^{1}$ is the successor state of $S_{0}^{0}$, yet we need to complete the search for $S_{0}^{1}$ first. Fortunately, this is possible during the reanalyze process.
@@
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order. Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$, and $S_{l}^{t}$ is the $t$-th state in the $l$-th trajectory. We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on. After we search on state $S_{l}^{t+1}$  [^2] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes., the root value $m_{l}^{t+1}$ is obtained. When engaging in search on $S_{l}^{t}$, we assign the value of $S_{l}^{t+1}$ to the fixed value $m_{l}^{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $S_{l}^{t}$ with the following equation:
@@
$I_{t}(a)=\begin{cases}UCB_{score}(S^{t}_{l},a),&a\neq a_{l}^{t}\\ r_{l}^{t}+\gamma m_{l}^{t+1},&a=a_{l}^{t}\end{cases}$ (5)
where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model. If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal in accordance with the original setting [10]. If action $a_{l}^{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version.
Figure 2: An illustration about the backward-view reanalyze in both node and batch view. We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories. From the node view, we would first search $S_{0}^{1}$ and then pass root value $m_{0}^{1}$ to $S_{0}^{0}$ to evaluate the value of a child node. $T_{0}^{1}$ and $T_{0}^{0}$ are the corresponding search trees. From the batch view, we would group all $S^{1}$s into a sub batch to search together and pass the root values to the $S^{0}$s.

### v5

@@ Abstract
Monte Carlo Tree Search (MCTS)-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains. These algorithms employ the reanalyze process to enhance sample efficiency from stale data, albeit at the expense of significant wall-clock time consumption. To address this issue, we propose a general approach named ReZero to boost tree search operations for MCTS-based algorithms. Specifically, drawing inspiration from the one-armed bandit model, we reanalyze training samples through a backward-view reuse technique which uses the value estimation of a certain child node to save the corresponding sub-tree search time. To further adapt to this design, we periodically reanalyze the entire buffer instead of frequently reanalyzing the mini-batch. The synergy of these two designs can significantly reduce the search cost and meanwhile guarantee or even improve performance, simplifying both data collecting and reanalyzing. Experiments conducted on Atari environments, DMControl suites and board games demonstrate that ReZero substantially improves training speed while maintaining high sample efficiency. The code is available as part of the LightZero MCTS benchmark at https://github.com/opendilab/LightZero.
@@ 1 Introduction
As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) (Sutton & Barto, 1988) has acquired achievements and applications across diverse fields, including interactive gaming (Vinyals et al., 2019), autonomous vehicles (Li et al., 2022), and natural language processing (Rafailov et al., 2023). Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency. These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real-world scenarios. In response to this problem, numerous model-based RL methods (Janner et al., 2019; Hafner et al., 2020) have emerged. These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
@@
Within this field, Monte Carlo Tree Search (MCTS) (Świechowski et al., 2023) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in a wide range of scenarios. Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep reinforcement learning with MCTS, achieving notable accomplishments that can beat top-level human players. While it can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc. All these MCTS-based algorithms made valuable contributions to the universal applicability of the MCTS+RL paradigm.
@@
However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state. Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments (Weng et al., 2022), further amplifying the speed disadvantage. On the other hand, during the reanalyze (Schrittwieser et al., 2021) process, in order to obtain higher-quality update targets, the latest models are used to re-run MCTS on the training mini-batch. The wall-clock time thus increases as a trade-off for high sample efficiency. The excessive cost has become a bottleneck hindering the further promotion of these algorithms.
@@
Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead. On the one hand, SpeedyZero (Mei et al., 2023) diminishes algorithms’ time overheads by deploying a parallel execution training pipeline; however, it demands additional computational resources. On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra demands. PTSAZero (Fu et al., 2023), for instance, compresses the search space via state abstraction, decreasing the time cost per search. In contrast, we aim to adopt a method that is orthogonal to both of the previous approaches. It does not require state space compression but directly reduces the search space through value estimation, and it does not introduce additional hardware overhead.
@@
In this paper, we introduce ReZero, a new approach/framework designed to boost the MCTS-based algorithms. Firstly, inspired by the one-armed bandit model (Lattimore & Szepesvári, 2020), we propose a backward-view reanalyze technique that proceeds in the reverse direction of the trajectories, utilizing previously searched root values to bypass the exploration of specific child nodes, thereby saving time. Additionally, we have proven the convergence of our search mechanism based on the non-stationary bandit model, i.e., the distribution of child node visits will concentrate on the optimal node. Secondly, to better adapt to our proposed backward-view reanalyze technique, we have devised a novel pipeline that concentrates MCTS calls within the reanalyze process and periodically reanalyzes the entire buffer after a fixed number of training iterations. This entire-buffer reanalyze not only reduces the number of MCTS calls but also better leverages the speed advantages of parallelization. Skipping the search of specific child nodes can be seen as a pruning operation, which is common in different tree search settings. Reanalyze is also a common module in MCTS-based algorithms. Therefore, our algorithm design is universal and can be easily applied to the MCTS-based algorithm family. In addition, it will not bring about any overhead in computation resources. Empirical experiments (Section 5) show that our approach yields good results in both single-agent discrete-action environment (Atari (Bellemare et al., 2013)), two-player board games (Silver et al., 2016), and continuous control suites (Tunyasuvunakool et al., 2020), greatly improving the training speed while maintaining or even improving sample efficiency. Ablation experiments explore the impact of different reanalyze frequencies and the acceleration effect of backward-view reuse on a single search. The main contributions of this paper can be summarized as follows:
@@
  • We design a method to speed up a single tree search by the backward-view reanalyze technique. Theoretical support for the convergence of our proposed method is also provided.
@@
  • We propose an efficient framework with the entire-buffer reanalyze mechanism that further reduces the number of MCTS and enhance its parallelization, boosting MCTS-based algorithms.
@@
  • We conduct experiments on diverse environments and investigate ReZero through ablations.
@@ 2.1 MCTS-based Algorithms
AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero (Schrittwieser et al., 2019) combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. EfficientZero (Ye et al., 2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Another works (Hubert et al., 2021; Antonoglou et al., 2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. MuZero Unplugged (Schrittwieser et al., 2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Despite the aforementioned excellent improvement techniques, the time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@ 2.2 MCTS Acceleration
Recent research has focused on accelerating MCTS-based algorithms. Mei et al. (2023) reduces the algorithm’s time overhead by designing a parallel system; however, this method requires more computational resources and involves some adjustments for large batch training. Fu et al. (2023) narrows the search space through state abstraction, which amalgamates redundant information, thereby reducing the time cost per search. KataGo (Wu, 2019) use a naive information reuse trick. They save sub-trees of the search tree and serve as initialization for the next search. However, our proposed backward-view reanalyze technique is fundamentally different from this naive forward-view reuse and can enhance search results while saving time. To our knowledge, we are the first to enhance MCTS by reusing information in a backward-view. Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which (Mei et al., 2023; Fu et al., 2023; Danihelka et al., 2022; Leurent & Maillard, 2020) are orthogonal to our contributions.
@@ 3.1 MuZero
MuZero (Schrittwieser et al., 2019) is a fundamental model-based RL algorithm that incorporates a value-equivalent model (Grimm et al., 2020a) and leverages MCTS for planning within a learned latent space. The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
@@
The representation model transforms last observation sequences $o_{t-l:t}$ into a corresponding latent state $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the next latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value $v_{t}$. These outputs are instrumental in guiding the agent’s action selection throughout its MCTS. Lastly the agent samples the best action $a_{t}$ following the searched visit count distribution. MuZero Reanalyze is an advanced version of the original MuZero. This variant enhances the model’s accuracy by conducting a fresh MCTS on sampled states with the latest model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection. Traditional uses of the algorithm have intertwined reanalysis with training, while we suggest a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology. Refer to the Appendix B for more details on MuZero during the training and inference phases.
@@ 3.2 Bandit-view Tree Search
A stochastic bandit has $K$ arms, and playing each arm means sampling a reward from the corresponding distribution. For a search tree in MCTS, the root node can be seen as a bandit, with each child node as an arm. The left side of Figure 1 illustrates this idea. However, as the policy is continuously improved during the search process, the reward distribution for the arms should change over time. Therefore, UCT (Kocsis & Szepesvári, 2006) modeled the root node as a non-stationary stochastic bandit with a drift condition:
$\mathbb{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}-\mu_{i}\geq\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})\>\text{and}\>\mathbb{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}-\mu_{i}\leq-\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})$ (2)
Where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}$ is the average reward of the first $s$ samples of arm $i$. $\mu_{i}$ is the limit of $\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}]$ as $s$ approaches infinity, which indicates that the expectation of the node value converges. $C$ is an appropriate constant characterizes the rate of concentration.
Based on this modeling, UCT uses the bandit algorithm UCB1 (Auer et al., 2002) to select child nodes. AlphaZero inherits this concept and employs a variant formula:
@@
where $s$ is the state corresponding to the current node, $a$ is the action corresponding to a child node. $Q(s,a)$ is the mean return of choosing action $a$, $P(s,a)$ is the prior score of action $a$, $N(s,a)$ is the total time that $a$ has been chosen, ${\sum}_{b}N(s,b)$ is the total time that $s$ has been visited. Viewing tree search from the bandit-view inspired us to use techniques from the field of bandits to improve the tree search. In next section, We use the idea of the one-armed bandit model to design our algorithm and prove the convergence of our algorithm based on the non-stationary bandit model.
@@ 4 Method
In this section, we introduce the specific design of ReZero. Section 4.1 describes the inspiration and specific operations of backward-view reanalyze. It performs a reverse search on the trajectory and uses the value estimation of a certain child node to save the corresponding sub-tree search time. Section 4.2 analyzes the node selection method introduced in Section 4.1 and ultimately proves the convergence of the method. Section 4.3 introduced the overall framework of ReZero, which adopts the buffer reanalyze setting to better integrate with the method described in Section 4.1.
@@ 4.1 Backward-view Reanalyze
Our algorithm design stems from a simple inspiration: if we could know the true state-value (e.g., expected long-term return) of a child node in advance, we could save the search for it, thus conserving search time. As shown on the right side of Figure 1, we directly use the true expectation $\mu_{A}$ to evaluate the quality of Arm A, thereby eliminating the need for the back-propagated value to calculate the score in Eq. 3. This results in the process occurring within the gray box in Figure 1 being omitted. Indeed, this situation can be well modeled as an one-armed bandit (Lattimore & Szepesvári, 2020), which also facilitates our subsequent theoretical analysis.
@@
Figure 1: The connection between MCTS and bandits. Left shows tree search in a bandit-view. When the action $A$ is selected, a return $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$. For the root node, the traversal, evaluation and back-propagation occurring in the sub-tree can be approximated as sampling from an non-stationary distribution. Thus it can be seen as a non-stationary bandit. Right shows the one-armed bandit case. Once the true value $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
@@
Driven by the aforementioned motivation, we aspire to obtain the expected return of a child node in advance. However, the true value is always unknown. Consequently, we resort to using the root value obtained from MCTS as an approximate substitute. Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$  [^1] The subscript denotes the trajectory, the superscript denotes the time step in the trajectory. in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$. Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$. However, there is a temporal contradiction. $S_{0}^{1}$ is the successor state of $S_{0}^{0}$, yet we need to complete the search for $S_{0}^{1}$ first. Fortunately, this is possible during the reanalyze process.
@@
During reanalyze, since the trajectories were already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order. Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$, and $S_{l}^{t}$ is the $t$-th state in the $l$-th trajectory. We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on. After we search on state $S_{l}^{t+1}$  [^2] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes., the root value $m_{l}^{t+1}$ is obtained. When engaging in search on $S_{l}^{t}$, we assign the value of $S_{l}^{t+1}$ to the fixed value $m_{l}^{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $S_{l}^{t}$ with the following equation:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{l}^{t}(a)}=\begin{cases}UCB_{score}(S^{t}_{l},a),&a\neq a_{l}^{t}\\ r_{l}^{t}+\gamma m_{l}^{t+1},&a=a_{l}^{t}\end{cases}$ (5)
where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model. If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal with the original setting as in MuZero. If action $a_{l}^{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version. Algorithm 1 shows the specific design with Python-like code. [^3] Our enhanced search process is implemented through $reuse\_MCTS()$, where $select\_root\_child()$ performs the action selection method of Equation 5. The $traverse()$ and $backpropagate()$ represent the forward search and backward propagation processes in standard MCTS.
Figure 2: An illustration about the backward-view reanalyze in node and batch view. We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories. From the node view, we would first search $S_{0}^{1}$ and then pass root value $m_{0}^{1}$ to $S_{0}^{0}$ to evaluate the value of a child node. $T_{0}^{1}$ and $T_{0}^{0}$ are the corresponding search trees. From the batch view, we would group all $S^{1}$s into a sub batch to search together and pass the root values to the $S^{0}$s.